# Dynamical stability of evolutionarily stable strategy in asymmetric games

### v2

@@ I Introduction
Evolutionarily stable strategy (ESS) is the most fundamental concept of evolutionary game theory Smith and Price (1973); Smith (1982); Thomas (1985), and its exploration has yielded a substantial body of scientific literature Smith (1982); Nowak (2006); Hofbauer and Sigmund (1998); Cressman (2003); Weibull (1997). In fact, one may remark Dawkins (1989) that this concept is only second to the Darwin’s contribution in advancing the evolutionary theory. No wonder that the paper Maynard Smith (1974) putting forward this concept has been included as one of eighty-eight foundational papers in complexity science Krakauer (2024) of the 20th century. In a large (mathematically, infinite) population with all individuals adopting a strategy (equivalently, a phenotype) that is evolutionarily stable, a small fraction of alternate strategy (called mutant strategy) cannot invade the population. Essentially, the concept of ESS can be understood as follows: ESS either renders a higher expected payoff through its performance against itself or, in case the strategy ties with that of the mutant’s, it fetches more expected payoff against the mutant compared to what the mutant would.
@@
The ESS was originally introduced for a population where all the individuals are identical (or have same role), from the strategic interaction point of view; in such a population, one says that symmetric games are played by the individuals. However, the concept of ESS in asymmetric games can be quite tricky. Asymmetric games refer to games that involve individuals categorized into more than one class—each with a distinct role—and the strategic interaction involves interaction between players in different roles. To complicate the matter further, in these games, players belonging to different roles may also have different strategy sets. If there is any interaction among players of same role, the interaction may be termed intra-specific interaction; whereas the interaction between players of different roles is aptly termed inter-specific. In this context, in this paper we follow the terminology that an asymmetric game may be called bimatrix games (sometimes also called truly asymmetric game) if only inter-specific interaction is present. The purpose of this paper is to critically discuss various extant definitions of ESS in bimatrix and asymmetric games. Note that henceforth, the phrase ‘asymmetric game’ will not include ‘bimatrix game’ which will be used as a separate phrase. Moreover, unless otherwise specified, we shall work assuming only two roles present in a population for simplicity and without much loss of generality.
@@
There have been multiple attempts to extend the concept of ESS to asymmetric and bimatrix games. Taylor Taylor (1979) defined asymmetric ESS as the resilient strategy that fetches more total average payoff—summation of average payoffs of every subpopulation of individuals with distinct role—than what a mutant strategy would do. Around the same time, Selten Selten (1980) and Schuster et al. Schuster et al. (1981) proposed that ESS for bimatrix games could be defined either in terms of total average payoff or more strictly, in terms of requiring that average payoffs in every role be more; interestingly, unlike the case of symmetric games, the ESS can not be a mixed strategy in bimatrix games Pohley and Thomas (1979); Schuster et al. (1981). Consequently, alternative definitions of ESS in asymmetric and bimatrix games were proposed to allow for the mixed ESS concept. For instance, Cressman Cressman (1992, 1996); Cressman and Tao (2014) weakens the aforementioned ESS condition by requiring that average payoff in any one of the role should be more than what corresponding mutant could get. This concept is termed two-species ESS—henceforth abbreviated as 2ESS. For bimatrix games, Hofbauer and Sigmund Hofbauer and Sigmund (1998) gave an—what we shall prove in this paper—equivalent definition of 2ESS, again with a view to allow for emergence of mixed ESS. The concept is termed Nash–Pareto pair.
@@
The push for getting mixed ESS in asymmetric and bimatrix games stems from the fact that the replicator equation Taylor and Jonker (1978); Cressman and Tao (2014)—the evolutionary dynamics in line with the Darwinian tenet of evolution—possesses asymptotically stable interior fixed points. It is naturally desirable that the game-theoretic concept of ESS correspond to such fixed points, in other words, the states of the population. Hence, we compare various definitions of ESS and find which one corresponds to asymptotic stability in which cases. Furthermore, we invoke the information-theoretic concept of relative entropy Kullback and Leibler (1951); Cover and Thomas (2005) and associate deeper meanings to the correspondences established.
@@ II Equivalence between various definitions of ESS in bimatrix Games
Consider a population consisting of two subpopulations denoted as X and Y. Let $\mathbf{x}=(x_{1},x_{2},\cdots,x_{m})$ represent a state of subpopulation X, where each $x_{i}$ represents the frequency of the $i$th phenotype in the subpopulation. Similarly, let $\mathbf{y}=(y_{1},y_{2},\cdots,y_{n})$ represent a state of subpopulation Y, where each $y_{j}$ represents the frequency of the $j$th type in the subpopulation. The state $(\mathbf{x},\mathbf{y})$ of the entire population is an element of simplotope, $\Sigma_{m}\times\Sigma_{n}$, which is a Cartesian product of two simplices of dimensions $m-1$ and $n-1$. In this context, $b_{ij}$ and $c_{ij}$, respectively, represent the payoff of type $i$ in X against type $j$ in Y and the payoff of type $i$ in Y against type $j$ in X. All such payoffs are collectively expressed through the $m\times n$ matrix ${\sf B}$ with elements $b_{ij}$’s and and $n\times m$ matrix ${\sf C}$ with elements $c_{ij}$’s.
Let $(\mathbf{x},\mathbf{y})$ denote any arbitrary state of the population. Suppose that $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ represents an evolutionarily stable state (ESS). As is the idea behind ESS Smith (1982), a state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ may be said to be an ESS if in the case of an emergence of a mutant state, say, $(\bf{x},\bf{y})$, on average an individual of host population is comparatively fitter than a mutant. Since practically mutations are rather rare, it makes sense for this criterion to hold good for infinitesimal mutant fraction in the context of infinitely large host population. Because the host population has more than one subpopulation, unlike what happens in the standard case of symmetric games, the idea of ESS can be expressed in various precise mathematical definitions.
@@ II.1 Mutants in both subpopulations
  1. Definition 1a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{x},\epsilon_{y})$, with $0<\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})~{}~{}~{}\text{\emph{and}}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (1)
@@
    (ii) and if $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$, then for all $\mathbf{{y}}\neq\mathbf{\hat{y}}$,
    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y};$ (3)
    (iii) and if $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$, then for all $\mathbf{{x}}\neq\mathbf{\hat{x}}$,
    $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (4)
    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}~{}\text{and}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (5)
$\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}~{}~{}\text{and}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}.$ (6)
Rather than comparing fitnesses of host individuals against the mutants in each subpopulation separately, one could compare combined fitness of host individuals in both the subpopulations with combined fitness of mutants in both the subpopulations Schuster et al. (1981); Selten (1980); Pohley and Thomas (1979); Taylor (1979). Accordingly, we write the following set of three equivalent definitions (aptly termed sum-definition):
[…]
    (ii) and $\forall(\mathbf{x},\mathbf{y})\in\Sigma_{m}\times\Sigma_{n}$ for which equality in (i) holds, then
@@
It is obvious that when Definitions 1a, 1b, and 1c hold, the corresponding Definitions 2a, 2b and 2c hold as well; the converse is not always true. We would like to remark that in line with analogous definitions in the case of symmetric games Hofbauer and Sigmund (1998), versions (a), (b) and (c) in the above two set of definitions may be called first principle definition, Maynard-Smith–Price definition and neighbourhood definition, respectively.
The equilibrium condition of Definition 2b (Eq. (8)) appears different from Eq. (6). So while strict NE obviously implies ESS under sum-definition, it is not obvious if the ESS implies NE always. To check this out, let’s write inequality (8) in the following form: $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}+(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}\geq 0$. Since this has to be true for all $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, if $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}<0$ for some $\mathbf{x}$, then one can always find a $\mathbf{y}$ such that $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}+(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}<0$ which means both the terms individually has to satisfy $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}\geq 0$ and $(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}\geq 0$ for all $({\bf x},{\bf y})$ so that inequality (8) holds good. Concluding, an ESS under sum-definition implies NE.
We emphasize that the definitions in this section require simultaneous appearance of mutants in both the subpopulations, i.e., $\epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$ as mentioned in Definitions 1a and 2a explicitly. In other words, states $(\mathbf{\hat{x}},\mathbf{y})$ and $(\mathbf{x},\mathbf{\hat{y}})$ are not mathematically allowed mutant states. This restriction is relaxed in the next definition as introduced in the next section.

### v3

@@ I Introduction
Evolutionarily stable strategy (ESS) is the most fundamental concept of evolutionary game theory Smith and Price (1973); Smith (1982); Thomas (1985), and its exploration has yielded a substantial body of scientific literature Smith (1982); Nowak (2006); Hofbauer and Sigmund (1998); Cressman (2003); Weibull (1997). In fact, one may remark Dawkins (1989) that this concept is only second to the Darwin’s contribution in advancing the evolutionary theory. No wonder that the paper Maynard Smith (1974) putting forward this concept has been included as one of eighty-eight foundational papers in complexity science Krakauer (2024) of the 20th century. In a large (mathematically, infinite) population with all individuals adopting a strategy (equivalently, a phenotype) that is evolutionarily stable, a small fraction of alternative strategy (called mutant strategy) cannot invade the population. Essentially, the concept can be understood as follows: ESS either renders a higher expected payoff through its performance against itself or, in case the strategy ties with that of the mutant’s, it fetches more expected payoff against the mutant compared to what the mutant would.
@@
The ESS was originally introduced for a population where all the individuals are identical (or have the same role), from the strategic interaction point of view; in such a population, one says that symmetric games are played by the individuals. However, the concept of ESS in asymmetric games can be quite tricky. Asymmetric games refer to games that involve individuals categorized into more than one class—each with a distinct role—and the strategic interaction involves interaction between players in different roles. To complicate the matter further, in these games, players belonging to different roles may also have different strategy sets. If there is any interaction among players of the same role, the interaction is termed intra-specific interaction; whereas the interaction between players of different roles is aptly termed inter-specific. In this context, in this paper we follow the terminology that an asymmetric game is called bimatrix games (sometimes also called truly asymmetric game) if only inter-specific interaction is present. The purpose of this paper is to critically discuss various extant definitions of ESS in bimatrix and asymmetric games. Note that henceforth, the phrase ‘asymmetric game’ will not include ‘bimatrix game’ which will be used as a separate phrase. Moreover, unless otherwise specified, we shall work assuming only two roles present in a population for simplicity and without much loss of generality.
@@
There have been multiple attempts to extend the concept of ESS to asymmetric and bimatrix games. Taylor Taylor (1979) defined asymmetric ESS as the resilient strategy that fetches more total average payoff—summation of average payoffs of every subpopulation of individuals with distinct role—than what a mutant strategy would do. Around the same time, Selten Selten (1980) and Schuster et al. Schuster et al. (1981) proposed that ESS for bimatrix games could be defined either in terms of total average payoff or more strictly, by requiring that the average payoffs in every role exceed those of potential mutants. Interestingly, unlike in the case of symmetric games, the ESS in bimatrix games cannot be a mixed strategy Pohley and Thomas (1979); Schuster et al. (1981). Consequently, alternative definitions of ESS in asymmetric and bimatrix games were proposed to allow for the mixed ESS concept. For instance, Cressman Cressman (1992, 1996); Cressman and Tao (2014) weakens the aforementioned ESS condition by requiring that the average payoff in at least one role must exceed what a corresponding mutant could achieve. This concept is termed two-species ESS—henceforth abbreviated as 2ESS. For bimatrix games, Hofbauer and Sigmund Hofbauer and Sigmund (1998) provided an alternative definition of ESS to allow for the emergence of mixed ESS. The concept is termed Nash–Pareto pair.
@@
The search for mixed ESS in asymmetric and bimatrix games is motivated by the fact that the replicator equation—an evolutionary dynamics model aligned with Darwinian principles—can exhibit asymptotically stable interior fixed points Taylor and Jonker (1978); Cressman and Tao (2014). It is naturally desirable that the game-theoretic concept of ESS correspond to such fixed points. Hence, we compare various definitions of ESS and find which one corresponds to asymptotic stability. Furthermore, we invoke the information-theoretic concept of relative entropy Kullback and Leibler (1951); Cover and Thomas (2005) and associate deeper meanings to the correspondences established.
@@ II Equivalence between various definitions of ESS in bimatrix Games
We begin our discussion of ESS within a population game where multiple phenotypes are present. Accordingly, it is natural to examine the frequency distribution of these phenotypes and to identify conditions under which a population state is evolutionarily stable. In other words, we seek criteria that define evolutionarily stable (population)-states. Importantly, there is a direct correspondence between an evolutionarily stable state and the mean population strategy being evolutionarily stable, that is, if a particular population state is evolutionarily stable, then the mean strategy of that population is also an evolutionarily stable strategy Cressman (1990, 1992). We explore this connection in detail in Sections III.1 and III.3 after introducing the various definitions of ESS. Henceforth, unless otherwise specified, the abbreviation ESS will refer exclusively to an evolutionarily stable state (ESS).
Consider a population consisting of two subpopulations denoted as X and Y. Let $\mathbf{x}=(x_{1},x_{2},\cdots,x_{m})$ represent a state of subpopulation X, where each $x_{i}$ represents the frequency of the $i$th phenotype in the subpopulation. Similarly, let $\mathbf{y}=(y_{1},y_{2},\cdots,y_{n})$ represent a state of subpopulation Y, where each $y_{j}$ represents the frequency of the $j$th type in the subpopulation. The state $(\mathbf{x},\mathbf{y})$ of the entire population is an element of simplotope, $\Sigma_{m}\times\Sigma_{n}$ [$\Sigma_{m}\equiv\{(x_{1},\cdots,x_{m})\in\mathbb{R}^{m}:\Sigma_{i=1}^{m}x_{i}=1\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ x_{i}\geq 0\}$], which is a Cartesian product of two simplices of dimensions $m-1$ and $n-1$. In this context, $b_{ij}$ and $c_{ij}$, respectively, represent the payoff of type $i$ in X against type $j$ in Y and the payoff of type $i$ in Y against type $j$ in X. All such payoffs are collectively expressed through the $m\times n$ matrix ${\sf B}$ with elements $b_{ij}$’s and and $n\times m$ matrix ${\sf C}$ with elements $c_{ij}$’s.
Let $(\mathbf{x},\mathbf{y})$ denote any arbitrary state of the population. Suppose that $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ represents an evolutionarily stable state (ESS). Following the idea of ESS Smith (1982), a state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ qualifies as an ESS if in the case of an emergence of a mutant state, say, $(\bf{x},\bf{y})$, on average an individual of host population is comparatively fitter than a mutant. Since practically mutations are rather rare, it makes sense for this criterion to hold good for infinitesimal mutant fraction in the context of infinitely large host population. Because the population under consideration has more than one subpopulations, unlike what happens in the standard case of symmetric games, the idea of ESS can be expressed in various precise mathematical definitions.
@@ II.1 Mutants in both subpopulations
  1. Definition 1a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\in\Sigma_{m}\times\Sigma_{n}$, such that $\mathbf{{x}}\neq\mathbf{\hat{x}}$ and $\mathbf{{y}}\neq\mathbf{\hat{y}}$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{x},\epsilon_{y})$, with $0<\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})\leavevmode\nobreak\ \text{\emph{and}}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (1)
    Here, $\epsilon_{x}$ and $\epsilon_{y}$ are the sufficiently small fractions of mutants inside the subpopulations X and Y respectively. Note that the constraints $\mathbf{x}\neq\mathbf{\hat{x}}$ and $\mathbf{y}\neq\mathbf{\hat{y}}$ on the mutant state are necessary—otherwise, no ESS can exist under Definition 1a. This follows trivially, as allowing all possible mutants would also include those of the form $(\mathbf{\hat{x}},\mathbf{y})$, which would always invalidate the first inequality in (1). This restriction on mutant states can also be imposed in a slightly different manner, leading to an alternative and-definition, as discussed in Appendix A.1.
  2. Definition 1b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})$, such that $\mathbf{{x}}\neq\mathbf{\hat{x}}$ and $\mathbf{{y}}\neq\mathbf{\hat{y}}$, the following conditions hold,
[…]
    (ii) and if $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$, then,
    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y};$ (3)
    (iii) and if $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$, then,
    $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (4)
  3. Definition 1c: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\mathbf{{x}}\neq\mathbf{\hat{x}}$ and $\mathbf{{y}}\neq\mathbf{\hat{y}}$
    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}\leavevmode\nobreak\ \text{\emph{and}}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (5)
$\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}.$ (6)
Rather than comparing fitnesses of host individuals against the mutants in each subpopulation separately, one could compare combined fitness of host individuals in both the subpopulations with combined fitness of mutants in both the subpopulations Schuster et al. (1981); Selten (1980); Pohley and Thomas (1979); Taylor (1979). Accordingly, we write the following set of three definitions (aptly termed sum-definition):
    Note that in this definition, one could relax the restriction—$\mathbf{x}\neq\mathbf{\hat{x}}$ and $\mathbf{y}\neq\mathbf{\hat{y}}$—on the mutant state, as explicitly incorporated in Definition 2a itself through adoption of the condition— $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$. In this case, there is no issue with allowing mutants of the form $(\mathbf{\hat{x}},\mathbf{y})$ or $(\mathbf{x},\mathbf{\hat{y}})$. For instance, for a mutant of the form $(\mathbf{\hat{x}},\mathbf{y})$, inequality (7) simplifies to $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}>\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$. This definition was originally introduced in Refs. Taylor (1979); Selten (1980); Schuster et al. (1981).
@@
    (ii) and $\forall(\mathbf{x},\mathbf{y})\in\Sigma_{m}\times\Sigma_{n}$ for which inequality in (i) holds as equality, then
@@
A few remarks are in order. We first mention that Definitions 2a, 2b, and 2c are equivalent if we assume $\epsilon_{x}=\epsilon_{y}=\epsilon$. Next, we would like to remark that in line with analogous definitions in the case of symmetric games Hofbauer and Sigmund (1998), versions (a), (b), and (c) in the above two set of definitions may be called first principle definition, Maynard-Smith–Price definition and neighbourhood definition, respectively.
The equilibrium condition of Definition 2b (Eq. (8)) appears different from Eq. (6). So while strict NE obviously implies ESS under sum-definition, it is not obvious that ESS always implies NE. To check this out, let us write inequality (8) in the following form: $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}+(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}\geq 0$. Since this has to be true for all $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, if $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}<0$ for some $\mathbf{x}$, then one can always find a $\mathbf{y}$ such that $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}+(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}<0$ which is a contradiction. This shows that both the terms individually has to satisfy $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}\geq 0$ and $(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}\geq 0$ for all $({\bf x},{\bf y})$ so that inequality (8) holds. Concluding, an ESS under sum-definition implies NE.
We emphasize that the definitions in this section require simultaneous appearance of mutants in both the subpopulations, i.e., $\epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$ as mentioned in Definitions 1a and 2a explicitly. This restriction is relaxed in the next definition as introduced in the next section.